2y=(x-5)squared+6
find: the center
the vertex's-
the foci points-
directrix -
and end point lactus-

Answer by stanbon(75887)

2y=(x-5)squared+6
----
Rearrange:
(x-5)^2 = 2(y-3)
4p = 2, so p = 1/2
It is a parabola opening upward.
---------------------------
find: the center------: It is a parabola. No center.
the vertex's----------: (5,3)
the foci points-------: (5,3+2) = (5,5)
directrix ------------: x = 5-2 = 3
and end point lactus--: ???
==================================
Cheers,
Stan H.